# Rua Dr. Bento Teobaldo Ferraz 271, 01140-070, São Paulo, SP, Brazil

### v2

@@ 10 $(N(N(36+3N)+81)-74)N(N-1)/120$ 89 226 481
Clearly, the degeneracies at each spin are a polynomial in $N$ of degree $\lfloor J/2\rfloor$ and must be expressible as an integer combination of dimensions of representations of $S_{N}$. While for any fixed $N$ this is not particularly helpful, for generic, sufficiently large $N$, we can use the polynomial dependence in $N$ to constrain the allowed representations. For example, for $J=4$ and $N$ generic we know from the explicit construction in Antunes:2022vtb that there is one singlet. Writing an ansatz in terms of representations whose dimensions are a polynomial of at most degree 2 in $N$,
$\frac{N(N-1)}{2}=n_{2}\,\textrm{dim}(N-2,2)+n_{11}\,\textrm{dim}(N-2,1,1)+n_{1}\,\textrm{dim}(N-1,1)+1$ (36)
has the unique integer solution $n_{1}=n_{2}=1\,,n_{11}=0$, which we also confirmed by explicitly constructing the operators. Similarly, it is easy to work out that for $J=5$ we have a single representation $n_{11}=1$ and that for $J=6$ we have $n_{3}=1,\,n_{2}=2,\,n_{11}=1,\,n_{1}=3$ along with 2 singlets. [^11] To argue this for $J=6$, we also need to know that $n_{111}=0$. This is clear because we can only have three indices for a spin-6 current if all of them are on $L_{-2}$ charges, the anti-symmetrization of which will give zero. For all $J\geq 7$, at least some of the non-singlet currents need to be constructed before there is enough information to conclude how many currents there are in other representations. However, already at spin $6$ we notice an interesting phenomenon for low $N$. The representation $(N-3,3)$ only exists for $N\geq 6$ and its dimension
@@
vanishes for $N=5$ and is negative for $N=4$. In fact this phenomenon becomes more and more frequent as we increase spin and start accessing representations of higher rank. Remarkably, at $J=6$ we find that the multiplicities for $N=5$ remain unchanged, while for $N=4$ the multiplicity $n_{2}=1$ is reduced by one. It would be interesting to understand whether identities of the type
@@
$J$ $d^{\prime}_{J,1}$ $d_{J,1}^{\prime(N=4)}$ $d_{J,1}^{\prime(N=5)}$ $d_{J,1}^{\prime(N=6)}$
@@ 9 $\binom{N}{4}(N(N(7N+137)+34)-168)(N-1)/24$ 326 3585 20250
These grow much faster and are polynomials in $N$ of degree $\lfloor J/2\rfloor+4$, which is already quite suggestive that currents have a very high chance of lifting, especially as $N$ increases. However, not only must the number of divergences match or exceed the number of currents, it must do so for each individual irreducible representation of $S_{N}$. By explicitly constructing all the primaries, we find that this is the case for all spins, irreducible representations and values $N$ except for a single case. For $J=3$ and $N=4$, we find that the multiplicities are $n_{0}=1,\,n_{1}=2,\,n_{11}=1$, accounting for the ten states. However, there are no divergences in the $(2,2)$ representation that are candidates to recombine with the corresponding spin 4 current! For the pure $\sigma$ deformation, this is a non-perturbative constraint ensuring the presence of additional currents. In fact, this is simply a global symmetry refined version of Zamolodchikov’s famous counting argument for the integrability of deformed minimal models Zamolodchikov:1987ti ; Zamolodchikov:1989hfa . [^12] This complements the approach of LeClair:1997gv which established integrability via a Toda field theory construction. See also section 5.2. However, for our infrared CFT, both $g_{\sigma}$ and $g_{\epsilon}$ are non-vanishing and therefore operators built on top of $\phi_{(1,3)}$ might recombine with the current. In ordinary conformal perturbation theory this would manifest itself at third order, with one $\epsilon$ and two $\sigma$ insertions.
@@ 3.3 Comments on finite $m$
A reasonable worry about the analysis of Antunes:2022vtb is whether the lifting of currents is expected to persist when the spins $J$ of the currents start scaling with some positive power of $m$. Indeed, conformal perturbation theory presumably breaks down when this happens. We claim that the physics at sufficiently large but finite $m$ should not be modified substantially when compared to the large $m$ expansion. Indeed, the fundamental mechanism that ensures the lifting of currents is the overwhelming growth of potential divergences with which they can recombine. As we showed in the subsection above, at fixed $J$ the degeneracies of divergences $d^{\prime}_{J,1}$ when compared to the degeneracies of currents $d_{J,0}$ scale as
$\frac{d^{\prime}_{J,1}}{d_{J,0}}\sim N^{4}\,,$ (39)
@@
$J$ $d_{J,0}(m=4)$ $d_{J,0}(m=\infty)$ $d_{J,1}^{\prime}(m=4)$ $d_{J,1}^{\prime}(m=\infty)$
@@
We see that the putative divergences continue to vastly outnumber the currents even for $J=m^{2}=16$. In fact, at large $N$ one observes that the reduction in the number of states is suppressed by at least $1/N^{4}$ compared to the overall number of states both for currents and divergences. While ideally we should refine this counting for each individual irreducible representation, we believe that the mechanism that ensures lifting of the currents should remain robust at large but finite $m$.
@@ 4 Broken conservation and the fate of the infrared
We will now explain how to turn the above results into a concrete algorithm for strengthening the evidence that coupled minimal model fixed points are irrational. This leads to the main result of this paper which is that currents of spin less than $10$ outside the $\widehat{\mathfrak{Vir}}$ identity multiplet all appear to lift. Due to the exceptional nature of $N=4$, this has been checked for $N=5,6,7$. The steps below will be similar to those that were used (and explained very tersely) in Antunes:2022vtb but without any assumptions about the currents or divergences being singlets of $S_{N}$.
@@ 4.1 Construction of operators
The previous section has given an algorithm to determine $d_{J,0}$ which is the dimension of the $(h,\bar{h})=(J,0)$ Verma module in the UV theory. To go further, it is important to actually build the currents in this Verma module so that their fate in the infrared can be determined by the aforementioned recombination technique. We have done this in a two step process — first by generating a set of currents that form a basis when $N$ is arbitrarily large and second by filtering the resulting list for finite values of $N$. At large $N$, the number of linearly independent currents for each irreducible representation and spin are given in Table 5.
@@
$n$ 1 2 3 8
Table 5: Degeneracies of currents refined by representation and spin for the irreducible representations of $S_{N}(N\gg 1)$ which appear for spin up to 10. These results agree with Table 2 and the singlet counting in Antunes:2022vtb as they must.
@@
for $N=4$. An important comment at this stage is that nothing in the algorithm we have outlined so far guarantees that it will produce irreps. Indeed, the element of (48) which is in the $(2,2)$ irrep is not $T_{3}$ but $T_{3}+3T_{1}$. When constructing primaries that have $c$ free indices and a prescribed index symmetry, the resulting basis for the kernel of $L_{1}^{1}+\dots+L_{1}^{N}$ and $L_{2}^{1}+\dots+L_{2}^{N}$ generally mixes primaries in all irreps compatible with this symmetry where the number of free indices may be less than $c$. One can determine the linear combinations that correspond to irreps by proceeding iteratively in $c$ and computing two-point functions but in practice we have chosen to simply verify the lifting of the currents returned by default. [^13] An aesthetically pleasing (but also computationally intensive) alternative would be projecting currents in the kernel onto irreducible representations after the fact. The projectors would have to be formed from invariant tensors of $S_{N}$ which can be counted with the techniques of bpr23 .
The second step is to make sure the set of currents we have generated remains linearly independent when the finiteness of $N$ is taken into account. Two things to do in this regard are quite obvious. Some representations of $S_{N}$ expressed in terms of $N$ simply do not exist for $N$ small enough so currents containing these components never need to be generated in the first place. As an example, the $(N-3,3)$ irrep first exists for $N=6$ leading to the spin-6 primary
@@
for $N=6$ and no equivalent for $N=4,5$. It is also possible to save some effort by never generating terms which multiply $a>N$ strings of Virasoro charges. These vanish because of our assumption that copy indices are all different but this phenomenon cannot occur below spin 10. The main way to remove redundant degrees of freedom, however, is simply by brute force. After generating a potentially overcomplete set of currents, which are stored in a format independent of $N$, it is possible to make the sums over $N$ values completely explicit so that the matrix of currents can be reduced. The runtime of this procedure does depend on $N$ but we have been able to carry it out for currents up to spin 10 without much difficulty. The result is that we have to remove
@@
for $N=4$.
@@
Once operators in the $(h,\bar{h})=(J,0)$ Verma modula are constructed, we also need to consider operators in $(h,\bar{h})=(J,1)$ built using four copies of $\phi\equiv\phi_{(1,2)}$. This introduces a few ingredients which were not present for the currents. In particular, when we loop the number of free indices in a given term from 00 to $c$, there is a further loop over how many of these free indices should appear on strings that include a copy of $\phi$. Whether $\phi$ is part of a string with a free index or the singlet needed to achieve a spin of $J-1$, it should not be acted upon by $L_{-2}$ since it has a null state at level 2. As already anticipated in Table 3, the number of divergence candidates can easily be in the thousands due to the number of admissible terms in the ansatz being similarly large. To apply $L_{1}^{1}+\dots L_{1}^{N}$ and $L_{2}^{1}+\dots L_{2}^{N}$ to this ansatz more quickly, it is possible to only account for index symmetrizations (obtained from deleting positions in a master symmetrization) at the very end. Given a single term like $T^{ijk}$, one can apply $L_{1}^{1}+\dots L_{1}^{N}$ and $L_{2}^{1}+\dots L_{2}^{N}$ and then add the necesssary orbits under index permutations instead of doing this first. Even with this optimization however, our code for some irreps of interest ran out of memory at $J=10$ thus forcing us to stop at $J=9$. To remove redundant degrees of freedom from the space of divergence candidates, we have only implemented the trivial methods for this — focusing on irreps which actually exist at finite $N$ and omitting terms with $a>N$ strings. The explicit row reduction needed to find a basis for the divergences is very slow and fortunately not necessary. The extra operators in our collection just do one of two things.

### v3

@@ 10 $(N(N(36+3N)+81)-74)N(N-1)/120$ 89 226 481
Clearly, the multiplicities must be expressible as a non-negative integer combination of dimensions of representations of $S_{N}$. It is of course very important to understand the refinement of the spectrum under global symmetry irreps. While the next subsection will introduce a refined partition function which achieves this decomposition, it is already possible to compute the multiplicities of irreps for some low spins in Table 2. For example, for $J=4$ and $N$ generic we know from the explicit construction in Antunes:2022vtb that there is one singlet. Writing an ansatz in terms of representations whose dimensions are a polynomial of at most degree 2 in $N$,
$\frac{N(N-1)}{2}=d^{(N-2,2)}_{4,0}\,\textrm{dim}(N-2,2)+d^{(N-2,1,1)}_{4,0}\,\textrm{dim}(N-2,1,1)+d^{(N-1,1)}_{4,0}\,\textrm{dim}(N-1,1)+1$ (37)
has the unique integer solution $d^{(N-1,1)}_{4,0}=d^{(N-2,2)}_{4,0}=1\,,d^{(N-2,1,1)}_{4,0}=0$.Similarly, it is easy to work out that for $J=5$ we have a single representation $d^{(N-2,1,1)}_{5,0}=1$ and that for $J=6$ we have $d^{(N-3,3)}_{6,0}=1,\,d^{(N-2,2)}_{6,0}=2,\,d^{(N-2,1,1)}_{6,0}=1,\,d^{(N-1,1)}_{6,0}=3$ along with 2 singlets. [^14] To argue this for $J=6$, we also need to know that $d^{(N-3,1,1,1)}_{6,0}=0$. This is clear because we can only have three indices for a spin-6 current if all of them are on $L_{-2}$ charges, the anti-symmetrization of which will give zero. For $J\geq 7$, the unrefined partition function is no longer enough. However, already at spin $6$ we notice an interesting phenomenon for low $N$. The representation $(N-3,3)$ only exists for $N\geq 6$ and its dimension
@@
vanishes for $N=5$ and is negative for $N=4$. In fact this phenomenon becomes more and more frequent as we increase spin and start accessing representations of higher rank. Remarkably, at $J=6$ we find that the multiplicities for $N=5$ remain unchanged, while for $N=4$ the multiplicity $d^{(2,2)}_{6,0}=1$ is reduced by one. It would be interesting to understand whether identities of the type
@@
$J$ $d^{\prime}_{J+1,1}$ $d_{J+1,1}^{\prime(N=4)}$ $d_{J+1,1}^{\prime(N=5)}$ $d_{J+1,1}^{\prime(N=6)}$
@@ 9 $\binom{N}{4}(N(N(7N+137)+34)-168)(N-1)/24$ 326 3585 20250
These grow much faster and are polynomials in $N$ of degree $\lfloor J/2\rfloor+4$, which is already quite suggestive that currents have a very high chance of lifting, especially as $N$ increases. However, not only must the number of divergences match or exceed the number of currents, it must do so for each individual irreducible representation of $S_{N}$. We will soon see that this is the case for all spins, irreducible representations and values $N$ except for a single case. For $J=3$ and $N=4$, we find that the multiplicities are $d^{\prime(4)}_{4,1}=1,\,d^{\prime(3,1)}_{4,1}=2,\,d^{\prime(2,1,1)}_{4,1}=1$, accounting for the ten states. However, there are no divergences in the $(2,2)$ representation that are candidates to recombine with the corresponding spin 4 current! For the pure $\sigma$ deformation, this is a non-perturbative constraint ensuring the presence of additional currents. In fact, this is simply a global symmetry refined version of Zamolodchikov’s famous counting argument for the integrability of deformed minimal models Zamolodchikov:1987ti ; Zamolodchikov:1989hfa . [^15] The possibility of $\phi_{(1,3)}$-type operators recombining with the current along the RG flow of the pure $\sigma$ deformation is forbidden at finite $m$ by dimensional analysis. This complements the approach of LeClair:1997gv which established integrability via a Toda field theory construction. See also section 5.2. However, for our infrared CFTs, both $g_{\sigma}$ and $g_{\epsilon}$ are non-vanishing meaning that operators built on top of $\phi_{(1,3)}$ might recombine with the current. In ordinary conformal perturbation theory this would manifest itself at third order, with one $\epsilon$ and two $\sigma$ insertions which should generically be non-vanishing. There is therefore no reason to expect that the current should remain conserved at the fixed point, even for this special case.
3.3 Refined partition functions
While in practice we will have to explicitly construct the currents and candidate divergences in each representation of $S_{N}$ to check if anomalous dimensions are non-zero, it is still valuable to perform the counting without having to construct them. To do this we define a partition function refined by the irreducible representation under which the operators transform
$Z_{\lambda}(\tau,\bar{\tau})=\textrm{Tr}_{\mathcal{H_{\lambda}}}q^{L_{0}-\frac{c}{24}}\bar{q}^{\bar{L}_{0}-\frac{c}{24}}=\sum_{h,\bar{h}}d_{h,\bar{h}}^{\lambda}q^{h-\frac{c}{24}}\bar{q}^{\bar{h}-\frac{c}{24}}\,,$ (40)
where $\mathcal{H}_{\lambda}$ is the Hilbert space of operators transforming in the irrep $\lambda$ [^16] This is well defined since the global symmetry commutes with the stress-tensor and the dilatation operator is therefore block diagonal with each block associated to an irrep $\lambda$. and $d_{h,\bar{h}}^{\lambda}$ are the sought-after degeneracies of operators with weights $(h,\bar{h})$ in irrep $\lambda$. We claim that the refined partition function can be computed through the following formula
$Z_{\lambda}(\tau,\bar{\tau})=\frac{1}{|S_{N}|}\sum_{g}\chi_{\lambda}(g)Z(\tau,\bar{\tau};g)=\frac{1}{|S_{N}|}\sum_{[g]}|[g]|\chi_{\lambda}([g])Z(\tau,\bar{\tau};[g])\,,$ (41)
where $|S_{N}|=N!$ is the order of the symmetric group $[g]$ denotes the conjugacy class of the group element $g$ and $|[g]|$ is the size of this class. Additionally we introduced the twisted partition function
$Z(\tau,\bar{\tau};g)\equiv\textrm{Tr}_{\mathcal{H}}(q^{L_{0}-\frac{c}{24}}\bar{q}^{\bar{L}_{0}-\frac{c}{24}}g)=\sum_{h,\bar{h},\lambda}d_{h,\bar{h}}^{\lambda}q^{h-\frac{c}{24}}\bar{q}^{\bar{h}-\frac{c}{24}}\chi_{\lambda}(g)\,,$ (42)
where $\chi_{\lambda}(g)$ are the $S_{N}$ characters of representation $\lambda$. The last equality follows from the definition of $\chi$ as a trace and the fact that the global symmetry commutes with the conformal algebra, by definition. The twisted partition functions are easy to compute directly in terms of their untwisted counterpart since the group element $g$ acts by permuting the copies. Indeed, for a conjugacy class $[g]$ labeled by a partition of $N$ encoded in the row lengths $\lambda_{i}$ of the associated Young tableau with $n$ rows we find [^17] See also loop2 for a derivation of the twisted partition functions in the case of the $q$-state Potts model, where the $S_{q}$ symmetry acts on the single-site Hilbert space on a lattice.
$Z(\tau,\bar{\tau};g)=Z(\tau,\bar{\tau};[g])=\prod_{i=1}^{n}Z(\lambda_{i}\,\tau,\lambda_{i}\,\bar{\tau})\,.$ (43)
For example, for a permutation $(12)(3)(4)$, labeled by the partition $(2,1,1)$ we have
$Z(\tau,\bar{\tau};[(12)(3)(4)])=Z(2\tau,2\bar{\tau})Z(\tau,\bar{\tau})^{2}\,.$ (44)
To show why (43) holds, we make explicit use of the factorized Hilbert space. Using $\left|i_{1},\dots,i_{N}\right>$ as a shorthand for $\left|h_{i_{1}},\bar{h}_{i_{1}},\dots,h_{i_{N}},\bar{h}_{i_{N}}\right>$, we have
$Z(\tau,\bar{\tau};g)$ $=\sum_{i_{1},\dots,i_{N}}\langle i_{1},\dots,i_{N}|q^{L_{0}^{1}+\dots+L_{0}^{N}-\frac{c}{24}}\bar{q}^{\bar{L}_{0}^{1}+\dots+\bar{L}_{0}^{N}-\frac{c}{24}}|g\cdot i_{1},\dots,g\cdot i_{N}\rangle$
$=\sum_{i_{1},\dots,i_{N}}q^{h_{i_{1}}+\dots+h_{i_{N}}-\frac{c}{24}}\bar{q}^{\bar{h}_{i_{1}}+\dots+\bar{h}_{i_{1}}-\frac{c}{24}}\delta_{i_{1},g\cdot i_{1}}\dots\delta_{i_{N},g\cdot i_{N}}$ (45)
$=\sum_{i_{1},\dots,i_{n}}q^{\lambda_{1}(h_{i_{1}}-\frac{c}{24})}\bar{q}^{\lambda_{1}(\bar{h}_{i_{1}}-\frac{c}{24})}\dots q^{\lambda_{n}(h_{i_{n}}-\frac{c}{24})}\bar{q}^{\lambda_{n}(\bar{h}_{i_{n}}-\frac{c}{24})}=\prod_{i=1}^{n}Z(\lambda_{i}\,\tau,\lambda_{i}\bar{\tau})\,.$
It is now straightforward to derive (41) using the character orthogonality formula
$\frac{1}{|S_{N}|}\sum_{[g]}|[g]|\chi_{\lambda}([g])\chi_{\lambda^{\prime}}([g])=\delta_{\lambda,\lambda^{\prime}}\,.$ (46)
Indeed, using the form (42) for the twisted partition function, we can multiply by a character $\chi_{\lambda^{\prime}}(g)$ and sum over all group elements. Then, using the orthogonality relation (46), we immediately find the explicit form for the refined partition function (40), proving the master formula (41).
As an application let us write down the refined partition functions and degeneracies for the case $N=4$. Degeneracies for higher values of $N$ are listed in Appendix A. First, we recall the character table of $S_{4}$ in Table 4.
$S_{4}$ [()] [(12)] [(12)(34)] [(123)] [(1234)]
Class Size 1 6 3 8 6
$\chi_{(4)}$ 1 1 1 1 1
$\chi_{(3,1)}$ 3 1 -1 0 -1
$\chi_{(2,2)}$ 2 0 2 -1 0
$\chi_{(2,1,1)}$ 3 -1 -1 0 1
$\chi_{(1,1,1,1)}$ 1 -1 1 1 -1
Table 4: Character table for $S_{4}$ with the conjugacy classes labeled by a representative. The identity element is denoted by empty parentheses.
Then, we make use of the master formula (41) and of (43) to derive the refined partition functions for the singlet (4) and symmetric (2,2) representations
$Z_{(4)}(\tau,\bar{\tau})$ $=\frac{1}{4!}\left[Z(\tau,\bar{\tau})^{4}+6Z(2\tau,2\bar{\tau})Z(\tau,\bar{\tau})^{2}+3Z(2\tau,2\bar{\tau})^{2}+8Z(3\tau,3\bar{\tau})Z(q,\bar{q})+6Z(4\tau,4\bar{\tau})\right]\,,$
$Z_{(2,2)}(\tau,\bar{\tau})$ $=\frac{1}{4!}\left[2Z(\tau,\bar{\tau})^{4}+6Z(2\tau,2\bar{\tau})^{2}-8Z(3\tau,3\bar{\tau})Z(\tau,\bar{\tau})\right]\,,$ (47)
with similar formulae for the remaining irreps. It is then straightforward to expand these into characters, obtaining the degeneracies in Table 5 for currents and Table 6 for divergence candidates below.
$J$ $d^{(4)}_{J,0}$ $d^{(3,1)}_{J,0}$ $d^{(2,2)}_{J,0}$ $d^{(2,1,1)}_{J,0}$ $d^{(1,1,1,1)}_{J,0}$
0 1 0 0 0 0
1 0 0 0 0 0
2 0 1 0 0 0
3 0 0 0 0 0
4 1 1 1 0 0
5 0 0 0 1 0
6 2 3 1 1 0
7 0 1 1 2 0
8 4 6 4 3 0
9 1 5 2 5 2
10 6 14 8 8 1
Table 5: List of degeneracies for all $N=4$ diagonal-Virasoro primary currents distinguished by irreducible representations in the tensor product theory. We count the number of irreps using the same notation as (37). The number of states is obtained by multiplying by appropriate representation dimensions.
$J$ $d^{\prime(4)}_{J+1,1}$ $d^{\prime(3,1)}_{J+1,1}$ $d^{\prime(2,2)}_{J+1,1}$ $d^{\prime(2,1,1)}_{J+1,1}$ $d^{\prime(1,1,1,1)}_{J+1,1}$
0 1 0 0 0 0
1 0 1 0 0 0
2 0 1 1 0 0
3 1 2 0 1 0
4 2 3 2 2 0
5 2 6 3 4 0
6 5 10 6 6 2
7 7 17 9 13 2
8 12 27 18 21 5
9 18 46 25 37 9
Table 6: List of degeneracies for all diagonal-Virasoro primary divergence candidates built from four $\phi_{(1,2)}$ operators distinguished by irreducible representations in the tensor product theory.
@@ 3.4 Comments on finite $m$
A reasonable worry about the analysis of Antunes:2022vtb is whether the lifting of currents is expected to persist when the spins $J$ of the currents start scaling with some positive power of $m$. Indeed, conformal perturbation theory presumably breaks down when this happens. We claim that the physics at sufficiently large but finite $m$ should not be modified substantially when compared to the large-$m$ expansion. Indeed, the fundamental mechanism that ensures the lifting of currents is the overwhelming growth of potential divergences with which they can recombine. As we showed in the subsection above, at fixed $J$ the degeneracies of divergences $d^{\prime}_{J,1}$ when compared to the degeneracies of currents $d_{J,0}$ scale as
$\frac{d^{\prime}_{J+1,1}}{d_{J,0}}\sim N^{4}\,,$ (48)
@@
$J$ $d_{J,0}(m=4)$ $d_{J,0}(m=\infty)$ $d_{J+1,1}^{\prime}(m=4)$ $d_{J+1,1}^{\prime}(m=\infty)$
@@
We see that the putative divergences continue to vastly outnumber the currents even for $J=m^{2}=16$. In fact, at large $N$ one observes that the reduction in the number of states is suppressed by at least $1/N^{4}$ compared to the overall number of states both for currents and divergences. This counting can be refined for each individual irreducible representation as detailed in section 3.3 and the mechanism that ensures lifting of the currents should therefore remain robust at large but finite $m$.
@@ 4 Broken conservation and the fate of the infrared
We will now explain how to turn the above results into a concrete algorithm for strengthening the evidence that coupled minimal model fixed points are irrational. This leads to the main result of this paper which is that currents of spin less than $10$ outside the $\widehat{\mathfrak{Vir}}$ identity multiplet all lift for $N=5,6,7$. The steps below will be similar to those that were used (and explained very tersely) in Antunes:2022vtb but without any assumptions about the currents or divergences being singlets of $S_{N}$.
@@ 4.1 Construction of operators
The previous section has given an algorithm to determine $d_{J,0}$ which is the dimension of the $(h,\bar{h})=(J,0)$ Verma module in the UV theory. Individual multiplicities of irreps $d^{\lambda}_{J,0}$, satisfying $\sum_{\lambda}d^{\lambda}_{J,0}\text{dim}\lambda=d_{J,0}$, were also calculated. To go further, it is important to actually build the currents in this Verma module so that their fate in the infrared can be determined by the aforementioned recombination technique. We have done this in a two step process — first by generating a set of currents that form a basis when $N$ is arbitrarily large and second by filtering the resulting list for finite values of $N$. At large $N$, the number of linearly independent currents for each irreducible representation and spin are given in Table 8.
@@
$d^{(N-3,2,1)}_{J,0}$ 1 2 3 8
Table 8: The degeneracies of spin $J\leq 10$ currents which are relevant for the first step of our algorithm for constructing operators. They have been refined by $S_{N}$ representation under the assumption that $N\gg 1$. These can be predicted by following the finite $N$ techniques of section 3 by raising $N$ until the numbers stabilize.
@@
for $N=4$. An important comment at this stage is that nothing in the algorithm we have outlined so far guarantees that it will produce irreps. Indeed, the element of (57) which is in the $(2,2)$ irrep is not $T_{3}$ but $T_{3}+3T_{1}$. When constructing primaries that have $c$ free indices and a prescribed index symmetry, the resulting basis for the kernel of $L_{1}^{1}+\dots+L_{1}^{N}$ and $L_{2}^{1}+\dots+L_{2}^{N}$ generally mixes primaries in all irreps compatible with this symmetry where the number of free indices may be less than $c$. The second step of our procedure, to be discussed next, makes it possible to determine the linear combinations that correspond to irreps. Given a set of UV currents and divergence candidates which form a basis for finite $N$, one can proceed iteratively in $c$ by computing two-point functions and reaching irreps by demanding orthogonality. The main roadblock here is that going from an infinite $N$ basis to a finite $N$ basis only appears to be practical for the currents. For this reason, we have chosen to simply verify the lifting of currents without sorting them into irreps. [^18] An aesthetically pleasing (but also computationally intensive) alternative would be projecting currents in the kernel onto irreducible representations after the fact. The projectors would have to be formed from invariant tensors of $S_{N}$ which can be counted with the techniques of bpr23 .
The second step, which we will perform for currents but not divergences, is about making sure that the operators we have generated remain linearly independent when the finiteness of $N$ is taken into account. Two things to do in this regard are quite obvious. Some representations of $S_{N}$ expressed in terms of $N$ simply do not exist for $N$ small enough so currents containing these components never need to be generated in the first place. As an example, the $(N-3,3)$ irrep first exists for $N=6$ leading to the spin-6 primary
@@
for $N=6$ and no equivalent for $N=4,5$. It is also possible to save some effort by never generating terms which multiply $a>N$ strings of Virasoro charges. These vanish because of our assumption that copy indices are all different but this phenomenon cannot occur below spin 10. These tricks only get us so far and we must ultimately take a more brute force approach to discarding currents which are not linearly independent. As with the discussion under (55), this involves representing the currents as rows in a matrix and then reducing it. What makes this matrix especially large is that it cannot be formed from currents which are stored in a format independent of $N$. The data structures for the currents need to be unpacked so that the sums over $N$ are completely explicit. As such, the runtime of this procedure does depend on $N$ but we have been able to carry it out for currents up to spin 10 without much difficulty. The result is that we have to remove
@@
for $N=4$, in keeping with the degeneracies in Appendix A. We can say more about the simplest example by considering the spin-6 currents with four copies. Those in the singlet representation $(4)$, first constructed in Antunes:2022vtb , are
$\begin{split}T_{5}&=\frac{97}{50}\Sigma L_{-6}+\frac{119}{25}\Sigma L_{-4}L_{-2}-\frac{1}{2}\Sigma L_{-2}\Sigma L_{-4}-\frac{343}{200}\Sigma L_{-3}L_{-3}\\ &+\frac{5}{16}\Sigma L_{-3}\Sigma L_{-3}-\frac{28}{25}\Sigma L_{-2}L_{-2}L_{-2}+\Sigma L_{-2}\Sigma L_{-2}\Sigma L_{-2}\\ T_{6}&=\frac{193}{50}\Sigma L_{-6}+\frac{236}{25}\Sigma L_{-4}L_{-2}-\frac{3}{2}\Sigma L_{-2}\Sigma L_{-4}-\frac{667}{200}\Sigma L_{-3}L_{-3}\\ &+\frac{9}{16}\Sigma L_{-3}\Sigma L_{-3}-\frac{57}{25}\Sigma L_{-2}L_{-2}L_{-2}+\Sigma L_{-2}\Sigma L_{-2}L_{-2}.\end{split}$ (63)
Adding currents in the standard representation $(3,1)$ to this list does not produce any overcompleteness but one can instead add
$T_{7}$ $=2L_{-3}^{i}L_{-3}^{j}+\frac{16}{9}L_{-2}^{i}L_{-2}^{j}(L_{-2}^{i}+L_{-2}^{j})-\frac{8}{3}(L_{-2}^{i}L_{-4}^{j}+L_{-4}^{i}L_{-2}^{j})$
$+\frac{4}{5}\Sigma L_{-2}(L_{-4}^{i}+L_{-4}^{j})-\Sigma L_{-3}(L_{-3}^{i}+L_{-3}^{j})-\frac{96}{5}\Sigma L_{-2}\Sigma L_{-2}(L_{-2}^{i}+L_{-2}^{j})$
$+\frac{16}{9}\Sigma L_{-2}L_{-2}(L_{-2}^{i}+L_{-2}^{j})-\frac{4}{15}\Sigma L_{-4}(L_{-2}^{i}+L_{-2}^{j})+\frac{9056}{1125}\Sigma L_{-2}L_{-2}L_{-2}-2\Sigma L_{-3}\Sigma L_{-3}$
$+\frac{42226}{3375}\Sigma L_{-3}L_{-3}+\frac{16}{5}\Sigma L_{-2}\Sigma L_{-4}-\frac{116464}{3375}\Sigma L_{-4}L_{-2}-\frac{47416}{3375}\Sigma L_{-6}$
$T_{8}$ $=2\Sigma L_{-2}L_{-2}^{i}L_{-2}^{j}-\Sigma L_{-2}\Sigma L_{-2}(L_{-2}^{i}+L_{-2}^{j})+\frac{28}{75}\Sigma L_{-2}L_{-2}L_{-2}-\frac{5}{48}\Sigma L_{-3}\Sigma L_{-3}$
$+\frac{343}{600}\Sigma L_{-3}L_{-3}+\frac{1}{6}\Sigma L_{-2}\Sigma L_{-4}-\frac{119}{75}\Sigma L_{-4}L_{-2}-\frac{97}{150}\Sigma L_{-6}$ (64)
which both appear to have a $(2,2)$ component in this notation. To check whether this is really true, we must appreciate the fact that the 8 types of terms written in (63) and the further 9 types of terms with free indices in (64) can only be regarded as independent abstract quantities when $N$ is sufficiently large. Since $N=4$ here, we must represent these operators as vectors in a space of dimension 50 instead of 17. The basis is comprised of 4 $L_{-6}^{i}$ terms, 10 $L_{-3}^{i}L_{-3}^{j}$ terms, 16 $L_{-2}^{i}L_{-4}^{j}$ terms and 20 $L_{-2}^{i}L_{-2}^{j}L_{-2}^{k}$ terms. Once $T_{5}$ and $T_{8}$ are written in this higher dimensional format, it becomes easy to see that $T_{8}$ always comes out to be $-\frac{1}{3}T_{5}$ regardless of which values are chosen for its free indices $i$ and $j$. We therefore verify the last line of (62) by noticing that one operator designed to be new was secretly proportional to a singlet generated in a previous step.
@@
Once operators in the $(h,\bar{h})=(J,0)$ Verma module have been constructed using the two step process above, we need to repeat at least the first step for operators in $(h,\bar{h})=(J+1,1)$ built using four copies of $\phi\equiv\phi_{(1,2)}$. This introduces a few ingredients which were not present for the currents. In particular, when we loop the number of free indices in a given term from 00 to $c$, there is a further loop over how many of these free indices should appear on strings that include a copy of $\phi$. Whether $\phi$ is part of a string with a free index or the singlet needed to achieve a spin of $J-1$, it should not be acted upon by $L_{-2}$ since it has a null state at level 2. As already anticipated in Table 3, the number of divergence candidates can easily be in the thousands due to the number of admissible terms in the ansatz being similarly large. To apply $L_{1}^{1}+\dots L_{1}^{N}$ and $L_{2}^{1}+\dots L_{2}^{N}$ to this ansatz more quickly, it is possible to only account for index symmetrizations (obtained from deleting positions in a master symmetrization) at the very end. Given a single term like $T^{ijk}$, one can apply $L_{1}^{1}+\dots L_{1}^{N}$ and $L_{2}^{1}+\dots L_{2}^{N}$ and then add the necesssary orbits under index permutations instead of doing this first. Even with this optimization however, our code for some irreps of interest ran out of memory at $J=10$ thus forcing us to stop at $J=9$. To remove redundant degrees of freedom from the space of divergence candidates, we have only implemented the trivial methods for this — focusing on irreps which actually exist at finite $N$ and omitting terms with $a>N$ strings. The explicit row reduction needed to find a basis for the divergences is very slow and fortunately not necessary. The extra operators in our collection just do one of two things.